A long copper wire with radius of 1.0 mm carries a uniform surface charge density of 5.0 x 10^-6 C/m².






a. Find the total charge in a 1.0-meter-long section of the wire.






b.. Find the magnitude of the electric field at a distance of 15 cm from the wire.

1
Expert's answer
2021-12-21T12:16:23-0500

a. The total charge is


"Q=qA=q(2\pi rL)=3.14·10^{-8}\text{ C}."


b. The electric field is


"E=\frac{Q/L}{2\pi\epsilon_0R}=3763\text{ V/m}."